Given the recurrence relation π‘Žπ‘› =βˆ’2π‘Žπ‘›βˆ’1 +15π‘Žπ‘›βˆ’2 with the initial conditions π‘Ž0=1 and π‘Ž1=7.





(a) Write the characteristic equation.





(b) Solve the recurrence relation.






1
Expert's answer
2022-04-06T16:33:59-0400

"a_n=-2a_{n-1}+15a_{n-2}\\\\a:\\\\\\lambda ^2=-2\\lambda +15\\\\\\lambda ^2+2\\lambda -15=0\\\\\\lambda \\in \\left\\{ -5,3 \\right\\} \\\\b:\\\\a_n=C_1\\left( -5 \\right) ^n+C_23^n\\\\\\left\\{ \\begin{array}{c}\ta_0=1\\\\\ta_1=7\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_1+C_2=1\\\\\t-5C_1+3C_2=7\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_1=-0.5\\\\\tC_2=1.5\\\\\\end{array} \\right. \\\\a_n=-0.5\\left( -5 \\right) ^n+1.5\\cdot 3^n"